An amount of sand is warmed up from 15 degrees celcius to 85 degrees celcius. the mass of the sand is 300 gram. Calculate the amount of energy needed for this. Show your working.
1
Expert's answer
2019-07-03T08:49:06-0400

We can find the amount of energy needed for warm up the sand from the formula:


"Q = mc \Delta T,"

here, "Q" is the amount of energy needed for warm up the sand, "m = 0.3 kg" is the mass of the sand, "c = 830 J/kg^{\circ}C" is the specific heat capacity of sand [1], "\Delta T = 85^{\circ}C - 15^{\circ}C = 70^{\circ}C" is the tempperature difference.

Then, from this formula we can calculate the amount of energy needed for warm up the sand:


"Q = mc \Delta T = 0.3 kg \cdot 830 \dfrac{J}{kg^{\circ}C} \cdot 70^{\circ}C = 1.74 \cdot 10^4 J."

Answer:

"Q = 1.74 \cdot 10^4 J."


[1] - https://www.engineeringtoolbox.com/specific-heat-capacity-d_391.html.